OpenStudy (anonymous):

Which expression is the simplified form of the square root of r to the thirty-third power? r^11 r^16 r^11sqrt(r). r^16sqrt(r).

OpenStudy (anonymous):

@ivettef365

OpenStudy (ivettef365):

that is \(\sqrt{r^33}\)

OpenStudy (ivettef365):

it is ^33

OpenStudy (anonymous):

r^33 yes

OpenStudy (ivettef365):

that is equal to \(\sqrt{r^16r^16r}\)

OpenStudy (anonymous):

ok

OpenStudy (ivettef365):

now you can take out one sqrt r^16, and the r by itself stays inside so you have r^16\(\sqrt{r}\)

OpenStudy (anonymous):

ok that makes sense! THANKS!!!!!!!!!!!!

OpenStudy (ivettef365):

yw :)